# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v3

@@ 3.1.1. Recursive Construction:
(40) $\bigoplus_{j=1}^{t}\left(C^{\prime}_{uj}\oplus{C^{\prime}}_{uj}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}$
and its last $s$ constituent codes are
@@
(41) $\left(\bigoplus_{i=1}^{s-1}\;C_{ui}\right)\oplus\mathcal{C}_{u-1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell},$
@@
where the $C_{ui}$ are HSO codes for all $1\leq i\leq s-1$ and $\mathcal{C}_{u-1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}\cong\mathbb{F}_{q}^{m^{u-1}\ell}$ is an ESO constituent code of $\mathcal{C}_{u}$.
In the result, we determine a formula for the dimension of $\mathcal{C}_{u}$ in our recursive construction. To compute the dimension in Lemma 3.1, we focus only on the special case of this construction where the HSO constituent codes of $\mathcal{C}_{u}$ are $m^{u-1}$-copies of the HSO constituent codes of $\mathcal{C}_{1}$, namely, $C_{ui}=m^{u-1}C_{1i}$, for all $1\leq i\leq s-1$. We invoke this condition to have some control on the dimension of each $\mathcal{C}_{u}$. We note that since each $C_{1i}$ are HSO, each of the $C_{ui}$ are also HSO by a generalization of Part (i) of Theorem 2.8. Moreover, if each $C_{1i}$ has parameters $\ell,k_{1i},d_{1i}]$, then the respective parameters for $C_{ui}$ are $m^{u-1}\ell,k_{1i},m^{u-1}d_{1i}]$ for all $1\leq i\leq s-1$.
@@
Let
$R_{q,m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq i\leq s$, with $f_{s}=x-1$. Then, for $u\geq 1$, there exists a $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with constituent codes (40) and (41), whose its dimension is
$\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}.$
where $k_{1i}=\dim(C_{1i})$ for all $1\leq i\leq s$.
Proof.
@@
(42) $\mathcal{C}_{1}\cong\left(\bigoplus_{j=1}^{t}\left({C_{1j}^{{}^{\prime}}}\oplus{C_{1j}^{{}^{\prime}}}^{\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}C_{1i}\right).$
Using Lemma 2.1, the fact that $\deg(g_{j}(x))=\deg(g_{j}^{*}(x))$ and $\dim({C^{\prime}}_{1j}^{\perp_{E}})=\ell-k^{\prime}_{1j}$ for all $1\leq j\leq t$, the dimension of $\mathcal{C}_{1}$ is
(43) $\dim(\mathcal{C}_{1})$ $=$ $\left(\sum_{j=1}^{t}\deg(g_{j})k_{1j}^{\prime}+\deg(g_{j}^{*})(\ell-k_{1j}^{\prime})\right)+\left(\sum_{i=1}^{s}\deg(f_{i})k_{1i}\right)$
$=$ $\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}.$
For any $u\geq 2$, let
(44) $\mathcal{C}_{u}=\bigoplus_{j=1}^{t}\left(C_{uj}^{\prime}\oplus{C^{\prime}}_{uj}^{\perp_{E}}\right)\oplus\left(\bigoplus_{i=1}^{s-1}C_{ui}\right)\oplus\mathcal{C}_{u-1},$
@@
where $C_{ui}$ is $m^{u-1}$-copies of $C_{1i}$, namely, $C_{ui}=m^{u-1}C_{ui}$, for all $1\leq i\leq s$. Given its recursive argument, $\dim(\mathcal{C}_{u})$ is computed via induction. For $u=1$, it is done. Now assume that the argument is valid for $u$, i.e., there exists a QC code $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ with the following dimension
(45) $\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
@@
where $k_{1i}=\dim(C_{1i})=\dim(m^{u-1}C_{1i})=\dim(C_{ui})$ for all $1\leq i\leq s-1$. Now let $\mathcal{C}_{u+1}\subset\mathbb{F}_{q}^{m^{u+1}\ell}$ be the QC code with the following $m^{u}\ell$-length constituent codes
@@
(46) $\mathcal{C}_{u+1}\cong\left(\bigoplus_{j=1}^{t}\left(C_{u+1,j}^{\prime}\oplus C_{u+1,j}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}{C_{u+1,i}}\right)\oplus\mathcal{C}_{u}$
where each $C_{u+1,i}=m^{u}C_{1i}$ are also HSO for all $1\leq i\leq s-1$. (See Theorem 2.8.) Using equation  (45), we have
@@
$\dim(\mathcal{C}_{u+1})$ $=\sum_{j=1}^{t}k^{\prime}_{u+1,j}\deg(g_{j})+\sum_{j=1}^{t}\deg(g_{j}^{*}){k^{\prime}}_{u+1,j}^{\perp_{E}}+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\dim(\mathcal{C}_{u})$
@@
$=\sum_{j=1}^{t}\left[\deg(g_{j})k^{\prime}_{u+1,j}+\deg(g_{j})\left(m^{u}\ell-k^{\prime}_{u+1,j}\right)\right]+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\dim(\mathcal{C}_{u})$
$=m^{u}\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\dim(\mathcal{C}_{u})$
$=m^{u}\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
$=\ell\left(m^{u}+\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+(u+1)\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
$=\ell\left(\frac{m^{u+1}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+(u+1)\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
as desired. ∎
In the following lemma, we utilize Theorem 2.9 to find a lower bound for the minimum distance $d(\mathcal{C}_{u})$ for each $u\geq 2$. To obtain our lower bound, we include the condition that for each $u\geq 2$, the last HSO constituent codes $C_{ui}$ in (41) satisfy
(47) $d(C_{ui})\geq md(C_{u-1,i}),$
@@
for all $1\leq i\leq s-1$. Note that repeated application of (47) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each $i$. This condition can be satisfied, for instance, by setting $C_{ui}=mC_{u-1,i}$. However, unlike in Lemma 3.1, taking copies of constituent codes is not required to obtain our result.
@@
Let
$R_{q,m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq i\leq s$ with $f_{s}=x-1$. Given the constituent codes provided in (40) and (41), then, for $u\geq 1$, there exists an $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$, where $d_{1s}=d(C_{1s})$.
Proof.
We again use the induction argument along with the associated cyclic codes per (22). We first note that $D_{2t+s}$ is the $m$-length repetition code, so $d(D_{2t+s})=m$. For $u=1$, equation (24) and the fact that all nonzero constituent codes of $\mathcal{C}_{1}$ satisfy (25) yields
$R_{2t+s}(\mathcal{C}_{1})$ $=$ $d(D_{2t+s})d(C_{1s})=md_{1s}\geq d_{1s}$
$R_{2t+s-1,2t+s}(\mathcal{C}_{1})$ $=$ $(d(C_{1,s-1})-d(C_{1s}))d(D_{2t+s-1})+d(C_{1s})d(D_{2t+s-1,2t+s})\geq d(C_{1s})=d_{1s}$

### v4

@@ 3.1 Recursive Construction of Infinite Families of QC Codes
$\bigoplus_{j=1}^{t}\left(C^{\prime}_{u,j}\oplus{C^{\prime}}_{u,j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}$ (27)
and whose last $s$ constituent codes are
@@
$\left(\bigoplus_{i=1}^{s-1}\;C_{u,i}\right)\oplus\mathcal{C}_{u-1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell},$ (28)
@@
where the $C_{u,i}$ are HSO codes for all $1\leq i\leq s-1$ and $\mathcal{C}_{u-1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}\cong\mathbb{F}_{q}^{m^{u-1}\ell}$ is an ESO constituent code of $\mathcal{C}_{u}$. So by Theorem 2.8,
$\mathcal{C}_{u}\cong\Bigg(\bigoplus_{j=1}^{t}\left(C_{u,j}^{\prime}\oplus{C^{\prime}}_{u,j}^{\perp_{E}}\right)\Bigg)\oplus\left(\bigoplus_{i=1}^{s-1}C_{u,i}\right)\oplus\mathcal{C}_{u-1},$ (29)
is an ESO $m^{u-1}\ell$-QC code of length $m^{u}\ell$ for all $u\geq 2$.
Remark 3.18.
Since the dual of an ESO code is EDC, our codes $\mathcal{C}_{u}^{\perp}$ will be an EDC $m^{u-1}\ell$-QC code of length $m^{u}\ell$ for all $u\geq 2$, in which case Proposition 2.10 is utilized.
3.2 The Dimension of the Family of Codes $\{\mathcal{C}_{u}\}_{u\geq 1}$
In the next lemma, we determine a formula for the dimension of each $\mathcal{C}_{u}$ in our recursive construction which depend exclusively on the dimensions of constituent codes of $\mathcal{C}_{1}$. To compute the dimension, we restrict to the special case in which the HSO constituent codes of $\mathcal{C}_{u}$ are obtained by taking $m^{u-1}$-fold copies of the HSO constituent codes of $\mathcal{C}_{1}$; that is, $C_{u,i}=m^{u-1}C_{1,i}$ for all $1\leq i\leq s-1$. This choice allows us to systematically control how the parameters of $\mathcal{C}_{u}$ evolve with $u$, since taking copies preserves the dimension while scaling the length and minimum distance in a predictable way; namely, since each $C_{1,i}$ is HSO, it follows from a generalization of Part (i) of Theorem 2.12 that each $C_{u,i}$ is also HSO. Moreover, if $C_{1,i}$ has parameters $\ell,k_{1,i},d_{1,i}]$, then $C_{u,i}$ has parameters $m^{u-1}\ell,\,k_{1,i},\,m^{u-1}d_{1,i}]$ for all $1\leq i\leq s-1$. While this assumption restricts us to a special case, it provides a tractable setting for analysis, and extensions to more general constructions will be explored in future work.
@@ 3.2 The Dimension of the Family of Codes $\{\mathcal{C}_{u}\}_{u\geq 1}$
Let $R_{q,m}^{\ell}$ be as in (25). Then for $u\geq 1$, there exists an $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with constituent codes (27) and (28), whose dimension is
$\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}.$
where $k_{1,i}=\dim(C_{1,i})$, for all $1\leq i\leq s$.
Proof 3.20.
@@
$\mathcal{C}_{1}\cong\left(\bigoplus_{j=1}^{t}\left({C_{1,j}^{{}^{\prime}}}\oplus{C_{1,j}^{{}^{\prime}}}^{\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}C_{1,i}\right).$
Using Lemma 2 and the facts that $\deg(g_{j})=\deg(g_{j}^{*})$ and $\dim({C^{\prime}}_{1,j}^{\perp_{E}})=\ell-k^{\prime}_{1,j}$ for all $1\leq j\leq t$, we have that
$\dim(\mathcal{C}_{1})$ $=$ $\left(\sum_{j=1}^{t}\deg(g_{j})k_{1,j}^{\prime}+\deg(g_{j}^{*})(\ell-k_{1,j}^{\prime})\right)+\left(\sum_{i=1}^{s}\deg(f_{i})k_{1,i}\right)$
$=$ $\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}.$
@@
For any $u\geq 2$, let $\mathcal{C}_{u}$ be the QC code defined in (29) where the $C_{u,i}$ are $m^{u-1}$-copies of $C_{1,i}$, namely, $C_{u,i}=m^{u-1}C_{u,i}$, for all $1\leq i\leq s$. Given its recursive argument, $\dim(\mathcal{C}_{u})$ is computed via induction. The case when $u=1$ is done. Now assume that the argument is valid for any integer $u$, i.e., there exists a QC code $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ with the following dimension
$\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}$ (30)
@@
where $k_{1,i}=\dim(C_{1,i})=\dim(m^{u-1}C_{1,i})=\dim(C_{u,i})$ for all $1\leq i\leq s-1$. Now let $\mathcal{C}_{u+1}\subset\mathbb{F}_{q}^{m^{u+1}\ell}$ be the QC code with the following $m^{u}\ell$-length constituent codes
@@
$\mathcal{C}_{u+1}\cong\left(\bigoplus_{j=1}^{t}\left(C_{u+1,j}^{\prime}\oplus C_{u+1,j}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}{C_{u+1,i}}\right)\oplus\mathcal{C}_{u}$
where each $C_{u+1,i}=m^{u}C_{1,i}$ are also HSO for all $1\leq i\leq s-1$ (see Theorem 2.12). Using (30), we have hus,
$dim_{q}\mathcal{C}_{u+1}$ $=$ $\sum_{j=1}^{t}\deg(g_{j})k_{uj}^{\prime}+(m^{u}\ell-\deg(g_{j}^{*})k_{uj}^{\prime})+\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+dim_{q}\mathcal{C}_{u}$ (31)
$=$ $\sum_{j=1}^{t}\deg(g_{j})k_{uj}^{\prime}+(m^{u}\ell-\deg(g_{j})k_{uj}^{\prime})+\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+\left(\frac{m^{u}-1}{m-1}\right)t\ell+k_{1s}+u\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}$
$=$ $m^{u}\ell t+\left(\frac{m^{u}-1}{m-1}\right)t\ell+(u+1)\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+k_{1s}$
$=$ $\left(\frac{m^{u+1}-1}{m-1}\right)t\ell+(u+1)\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+k_{1s}.$
$\dim(\mathcal{C}_{u+1})$ $=\sum_{j=1}^{t}k^{\prime}_{u+1,j}\deg(g_{j})+\sum_{j=1}^{t}\deg(g_{j}^{*}){k^{\prime}}_{u+1,j}^{\perp_{E}}+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+\dim(\mathcal{C}_{u})$
$=\sum_{j=1}^{t}\left[\deg(g_{j})k^{\prime}_{u+1,j}+\deg(g_{j})\left(m^{u}\ell-k^{\prime}_{u+1,j}\right)\right]+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+\dim(\mathcal{C}_{u})$
$=m^{u}\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}$
$=\ell\left(\frac{m^{u+1}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+(u+1)\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}$
as desired. $\qed$
3.3 A Lower Bound for the Minimum Distance of the Family of Codes $\{\mathcal{C}_{u}\}_{u\geq 1}$
In this section, we will use Jensen’s bound to find a lower bound for the minimum distance of the QC code $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ for each $u\geq 1$. o do so, we will use Jensen’s concatenated description for each $u\geq 2$,
@@ 3.3 A Lower Bound for the Minimum Distance of the Family of Codes $\{\mathcal{C}_{u}\}_{u\geq 1}$
$\mathcal{C}_{u}\cong\left(\bigoplus_{j=1}^{t}\left(D_{uj}^{\prime}\square C_{uj}^{\prime}\oplus D_{uj}^{{}^{\prime}\perp_{E}}\square C_{uj}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}D_{ui}\square C_{ui}\right)\oplus D_{u-1}\square\mathcal{C}_{u-1}.$ (32)
Let $\mathcal{C}_{u}$ be a ESO $m^{u-1}\ell$-QC code in $\mathbb{F}_{q}^{m^{u}\ell}$
@@
$\mathcal{C}_{u}\cong\left(\bigoplus_{j=1}^{t}\left(D_{uj}^{\prime}\square C_{uj}^{\prime}\oplus D_{uj}^{{}^{\prime}\perp_{E}}\square C_{uj}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}D_{ui}\square C_{ui}\right)\oplus D_{u-1}\square\mathcal{C}_{u-1}.$ (33)
where $C_{ui}$ is HSO for each $1\leq i\leq s-1$.
where each $D^{\prime}_{uj},D^{{}^{\prime}\perp_{E}}_{uj}$, and $D_{ui}$ are the cyclic subcodes of $D_{u}$ associated to $C^{\prime}_{uj},C^{{}^{\prime}\perp_{E}}_{uj}$, and $C_{ui}$ respectively, for $1\leq j\leq t$ for $1\leq i\leq s-1$, and $D_{u-1}$ is the cyclic code associated to $\mathcal{C}_{u-1}$. We will then utilize Jensen’s Bound (Theorem 4) to find a lower bound for the minimum distance $d(\mathcal{C}_{u})$ for each $u\geq 2$. In order to use Theorem 4 in conjunction with our recursive construction of $\mathcal{C}_{u}$ outlined in (27) and (28), the constituent codes of each $\mathcal{C}_{u}$ must satisfy the descending chain condition in (10). We first decompose $\mathcal{C}_{u}$ for $u\geq 2$ as
$\mathcal{C}_{u}\cong\bigoplus_{\lambda=1}^{2t+s}C_{u,\lambda}$
where the $C_{u,\lambda}$ are the nontrivial constituent codes of $\mathcal{C}_{u}$ whose minimum distances satisfy (10), noting that $\mathcal{C}_{u-1}=C_{u,\lambda}$ for some $\lambda$. In general, we do not have enough information to conclude how $d(\mathcal{C}_{u-1})$ compares with the minimum distances of the other constituent codes for $\mathcal{C}_{u}$ and do not know where it appears in the descending chain. Thus, we require the additional condition that the constituent code with the smallest minimum distance in $\mathcal{C}_{u}$ grows by a factor of $m$ from the smallest constituent code in $\mathcal{C}_{u-1}$, i.e.,
$d(\mathcal{C}_{u,2t+s})\geq md(\mathcal{C}_{u-1,2t+s}).$ (34)
This is not a major assumption as $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ and $\mathcal{C}_{u-1}\subset\mathbb{F}_{q}^{m^{u-1}\ell}$, so its minimum distance can be a factor of $m$ larger than $\mathcal{C}_{u-1}$. Repeated application of (34) yields
$d(\mathcal{C}_{u,2t+s})\geq m^{u-1}d_{1,s}$ (35)
@@
where $d_{1,s}=d(C_{1,s})$ in (26). We emphasize that having such a condition will be vital in proving a lower bound for the minimum distances of $\mathcal{C}_{u}$ for each $u$ in the following lemma. This condition can be satisfied, for instance, by setting $C_{u,\lambda}=mC_{u-1,\lambda}$ for each $u\geq 2$ and $1\leq\lambda\leq 2t+s$. However, unlike in Lemma 3.19, taking $m$ copies of constituent codes is not required to obtain our result.
@@
Let $R_{q,m}^{\ell}$ be as in (25). Then for $u\geq 1$, there exists an $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with constituent codes (27) and (28), with minimum distance $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,s}$, where $d_{1,s}=d(C_{1,s})$.
Proof 3.22.
The case for $\mathcal{C}_{1}$ is trivial as we are assuming that $C_{1,s}$ has the smallest minimum distance in the representation of $\mathcal{C}_{1}$ in (26). For each $u\geq 2$, we will assume that the constituent codes in the decomposition of $\mathcal{C}_{u}$ are ordered according to the descending chain condition (10). The proof will require that we traverse our table horizontally by way of the descending chain condition to the last constituent then up the table vertically to use the inequality (35). For each $u\geq 2$, using the descending chain condition (10) and inequality (35), we have the following:
$N_{2t+s}(\mathcal{C}_{u})$ $=d(C_{u,2t+s})d(D_{2t+s})\geq md(C_{u-1,2t+s})\geq\ldots\geq m^{u-1}d(C_{1,s})=m^{u-1}d_{1,s},$
$N_{2t+s-1,2t+s}(\mathcal{C}_{u})$ $=d(C_{u,2t+s-1})d(D_{2t+s-1,2t+s})\geq d(C_{u,2t+s})\geq m^{u-1}d(C_{1,s})=m^{u-1}d_{1,s},$